# Deepfake Detection that Generalizes Across Benchmarks

### v2

@@ Abstract
This work delivers a computationally efficient and reproducible method, proving that state-of-the-art generalization is attainable by making targeted, minimal changes to a pre-trained foundational image encoder model. The code will be made publicly available upon acceptance.
@@ 1 Introduction
The proliferation of realistic facial deepfakes raises significant concerns regarding misinformation and malicious use, with AI-manipulated videos – those altered by techniques like face swapping or face reenactment – making detection challenging. Unlike fully synthetic content, such forgeries preserve the original context and leave subtle artifacts that are difficult for humans and machines to detect [24, 10].
@@
The generalization gap is the primary issue that we address in this work. Assuming the hypothesis that adapted large-scale, pre-trained foundational vision encoder can serve as a general foundation for deepfake detection [40], we build the proposed method in three variants, using Contrastive Language-Image Pre-training (CLIP) [43], Perception Encoder (PE) [8] and DINO [46] models as feature extractors, which are known for their generalizable visual representations.
The proposed method consists of a vision encoder, whose outputs are L2-normalized. We then fine-tune only the parameters of the Layer Normalization blocks [42] while keeping the rest frozen. Additionally, we propose using metric learning in this L2 space to enhance generalization.
@@ 2 Related Work
Deepfake techniques have rapidly evolved. Key methods include: Face-swap [4], Face-reenactment [9], LipSync [41], and full face video synthesis [48, 14]. The first three approaches manipulate only small, localized areas of the video, typically the facial region, while leaving the rest of the footage untouched, making detection challenging. The techniques are continually improving, leveraging advanced models to produce increasingly seamless and realistic results without obvious, visually perceptible artifacts.
As deepfake production techniques advance, detection methods are also evolving to keep pace. The literature encompasses a taxonomy of detection strategies, which can be divided into two main branches: content-based and signal-based approaches. Content-based methods focus on visible or interpretable inconsistencies, utilizing inductive biases such as apparent blending boundaries [45], unsynchronized audio-visual movements [20], or motion artifacts [1]. Signal-based methods, on the other hand, detect subtle and often invisible traces left in the visual signal, fingerprints that reveal synthetic origins even when no overt artifacts are present. Signal-based techniques have become increasingly accurate, especially as newer deepfakes leave fewer perceptible cues. However, a persistent challenge remains: generalizing to previously unseen manipulation methods.
Recent efforts in detecting general AI-generated content [40, 29, 39, 36] and facial deepfakes specifically [18, 37, 22, 13, 58] have begun to adopt CLIP as a backbone for generalizable deepfake detection. These approaches can be broadly categorized by the type of model adaptation. Some methods, like Forensics Adapter [13] (ForAda), introduce architectural changes by adding a separate parallel network explicitly trained to identify artifacts left by blending. Other methods operate on the parameter space; for instance, Effort [58] uses Singular Value Decomposition (SVD) to decompose the model’s weights into orthogonal subspaces, freezing the principal components to preserve pre-trained knowledge while fine-tuning the residual components on forgery patterns.
In contrast to these approaches, our work proposes to modify a subset of parameters. LN-Tuning [42] adjusts the affine parameters of Layer Normalization blocks. The efficiency of LN-tuning in the general context was analyzed in [19, 51]. We demonstrate that it achieves competitive or state-of-the-art performance in the deepfake detection task.
As a complimentary approach to the passive detection methods of AI manipulations, there exist active defense approaches that protect real images by incorporating invisible watermarks [65, 3, 6, 62, 2, 61].
@@ 3.1 Model
Feature extractor. We use three different pre-trained image encoders: 1. CLIP ViT-L/14, trained by OpenAI [43]; 2. $\text{PE}_{\text{core}}\text{L}$, trained by Meta AI [8]; 3. DINOv3 ViT-L/16 [46], trained by Meta AI. After the encoder processes the image, we take only the $1024$-dimensional classification token and discard the rest.
Optimized parameters. The L2-normalized classification token is used as the input to the linear binary classifier, which gives logits for fake and real classes. In addition to the weights $W\in\mathbb{R}^{1024\times 2}$ and biases $b\in\mathbb{R}^{2}$ of the classification layer, we optimize the parameters of all Layer Normalization blocks [42, 19, 51]. This roughly constitutes 0.03% of the total model parameters; the remaining parameters of the encoder are frozen.
Loss function. The network is trained using a weighted combination of cross-entropy, uniformity, and alignment losses [52]
@@
Training algorithm. To optimize the model parameters, we use Adam optimizer [30] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ without weight decay. The weight precision is set to bfloat16. The learning rate is scheduled using a cosine cyclic rule [47]. Each cycle starts with a linear warm-up for one epoch, increasing from $10^{-5}$ to $3\times 10^{-4}$, and then decays using cosine scheduling over nine epochs to $10^{-5}$. For most runs, we observed that the training did not improve after two cycles (or 20 epochs). The batch size is set to 128 samples. For CLIP ViT-L/14, the loss coefficients of Eq. 1 are set to $\alpha=0.1$ and $\beta=0.5$, found on the validation set.
@@ 3.2 Data
Data preprocessing. We follow a standard practice of dataset preprocessing proposed in DeepfakeBench [56], which consists of the following steps:
@@
    Extract the largest face using RetinaFace [15].
@@
Training data. Following the standard protocol, we train the model on the FaceForensics++ dataset [44] (FF++), which consists of $3600$ videos, of which $720$ are real videos, and the rest, $4\times 720$, come from four different deepfake forgery methods. After preprocessing, we obtain approximately 115115k frames. The dataset was released in three different compression levels. Following all of the preceding work, we use the c23 compression version of the dataset. We augment the training set using image augmentations such as random horizontal flipping, random affine transformations, Gaussian blurring, color jitter, and JPEG compression, all implemented in the torchvision package. In ablations, we also experimented with changing the training set to FFIW [67] and DSv2 [5].
@@
Validation data. We noticed that the FF++ validation set is very similar to the training set. That makes the comparison of different training experiments on such a validation ineffective. Each experiment achieves the best validation AUROC quickly, shows almost no signs of overfitting, and does not provide any estimate of cross-dataset generalization. The FF++ validation set does not provide crucial information on how the methods will perform when trained for an extended period. Therefore, we create a custom validation set, which consists of data from validation or training splits of other datasets, namely CDFv3 [35], DSv1 and DSv2 [5], and FFIW [67]. We ensure that test samples from these datasets are not present in the custom validation set.
@@ 4 Experiments
2019 FF++ [44] 4 140 560
2019 DF [44] 1 140 140
2019 F2F [44] 1 140 140
2019 FS [44] 1 140 140
2019 NT [44] 1 140 140
2019 DFD [17] 5 363 3068 -2
2019 UADFV [60] 1 49 49
2019 DFDC [16] 8 2500 -185 2500 -111
2020 FSh [32] 1 140 140
2020 CDFv2 [34] 1 178 340
2021 FFIW [67] 3 1738 -3 1738 -3
2021 KoDF [31] * 6 403 1106
2021 FAVC [28] 4 500 20566 -22
2022 DFDM [27] 5 590 -2 1720 -2
2024 PGF [25] 10 762 13605
2024 IDF [54] * 9 18834 $23$
2024 DSv1.1 [5] 5 1416 1497
2025 DSv2 [5] 6 1863 1416
2025 CDFv3 [35] 22 178 2540 -1
@@ 4.1 Test benchmarks
For reporting the performance of the model in cross-dataset settings, we are using Celeb-DF-v2 (CDFv2) [34], Celeb-DF++ (CDFv3) [35], DeepFake Detection Challenge (DFDC) [16], Google’s DFD dataset [17], Face Forensics in the Wild (FFIW) [67], DeepSpeak v1.1 (DSv1.1) and DeepSpeak v2.0 (DSv2) [5], Korean DeepFake Detection Dataset (KoDF) [31], FakeAVCeleb (FAVC) [28], DeepFakes from Different Models (DFDM) [27], PolyGlotFake (PGF) [25], and IDForge (IDF) [54]. The statistics for the test part of the datasets are shown in  Tab. 1.
@@ 4.2 Cross-dataset evaluation
LipForensics [20] 2021 CVPR Video ResNet-18 82.4 73.5
@@
RealForensics [21] 2022 CVPR Video Modified CSN 86.9 75.9
SBI [45] 2022 CVPR Frame EFNB4 93.2 82.7 72.4 84.8
AUNet [64] 2023 CVPR Video Xception+ART 92.8 99.2 73.8 81.5
StyleDFD [12] 2024 CVPR Video 3D ResNet-50 89.0 96.1
@@
LAA-Net [38] 2024 CVPR Frame EFNB4 95.4 98.4 86.9
AltFreezing [53] 2024 CVPR Video 3D ResNet-50 89.5 98.5
NACO [63] 2024 ECCV Video ViT-B/16 89.5 76.7
RAE [50] 2024 ECCV Frame ViT-B/16 95.5 99.0 80.2
TALL++ [55] 2024 IJCV Video Swin-B 92.0 78.5
ProDet [11] 2024 NeurIPS Frame EFNB4 92.5 77.0
UDD [18] 2025 AAAI Frame CLIP ViT-B/16 93.1 95.5 81.2
P&P [59] 2025 CVPR Video CLIP ViT-L/14 94.7 96.5 84.3 92.1
DFD-FCG [22] 2025 CVPR Video CLIP ViT-L/14 95.0 81.8
ForAda [13] 2025 CVPR Frame CLIP ViT-L/14 95.7 97.2 87.2
Effort [58] 2025 ICML Frame CLIP ViT-L/14 95.6 96.5 84.3 92.1
@@
The proposed method achieves these results simply by averaging individual frame softmax probabilities, without complex post-processing or architectural modifications as described in Sec. 3.1. This contrasts with many contemporary methods that rely on complex multimodal analyses, such as modeling temporal inconsistencies between video frames or detecting audio-video discrepancies [20, 37, 21, 66, 22, 59]. The performance of the proposed method suggests that the features of the visual encoder, when properly tuned, are highly effective for this detection task.
We also evaluated the generalization of the proposed method against current SOTA models across a wide range of benchmark datasets. For this experiment, we selected the most recent open-source SOTA models that have made their code and pre-trained weights publicly available, namely ForAda [13] and Effort [58]. To ensure that each model was used as intended by its original authors, we utilized their officially released weights and integrated their data preprocessing pipelines which were used during model training. After reproducing the results on the test datasets that were used in these papers, we obtained the same or very similar results, presented in Tab. 5. The in-distribution results on FF++ are shown in Tab. 3.
@@
ForAda [13] 99.7 97.0 98.6 91.9 96.8
Effort [58] 99.4 93.2 98.4 84.6 93.9
@@
ForAda [13] 99.4 97.2 87.3 82.0 95.7 90.6 88.2 93.1 97.1 86.6 90.8 81.8 72.8 75.6 88.4
Effort [58] 97.4 95.2 85.4 91.2 93.2 92.5 88.1 92.4 98.2 84.9 96.0 82.1 64.4 78.7 88.5
@@
ForAda [13] 95.7 / 95.7 97.2 / 97.2 87.2 / 87.3 – / 90.6
Effort [58] 95.6 / 93.2 96.5 / 95.2 84.3 / 85.4 92.1 / 92.5
@@ 4.3 Ablation studies
    Baseline: the setup is similar to [40]. We freeze the backbone and train a linear classifier with cross-entropy loss on top of the feature space of the classification token.
@@
We explore further parameter-efficient fine-tuning (PEFT) strategies, the goal of which is to modify only a relatively small number of parameters while keeping most parameters untouched, thus efficiently introducing several degrees of freedom. PEFT is more effective for generalization than full fine-tuning (FFT) in low-data regimes [23], which is particularly applicable to deepfake detection, where collected datasets and benchmarks have a limited number of samples. This work explores three PEFT methods: LoRA [26], LN-tuning [42, 33], and bias tuning (BitFit) [33, 7].
We found that FFT leads to rapid overfitting and poor generalization across datasets. This observation is consistent with recent findings from Effort [58] that FFT degrades the generalization performance.
Similarly to [58], we fine-tuned the parameters of every transformer weight matrix using LoRA to preserve the generalization. Setting LoRA to rank one, the lowest possible, allowed the model to reach a near-perfect training AUROC (99.99%) in two epochs, resulting in rapid overfitting.
We observed that BitFit [7] and LN-tuning [42] have stable training dynamics by first improving the validation AUROC, reaching a peak, and overfitting after some epochs. In comparison, FFT and LoRA show no improvement in validation AUROC even after a single training epoch.
@@
Uniformity and Alignment. We hypothesized that representations that effectively exploit the latent space could help us improve generalization. We used uniformity and alignment loss [52], leading to slower overfitting and better validation AUROC. We suppose that adding these losses helps to utilize the hyperspherical space more effectively, resulting in improved generalization.

### v3

@@ Abstract
This work delivers a computationally efficient and reproducible method, proving that state-of-the-art generalization is attainable by making targeted, minimal changes to a pre-trained foundational image encoder model. The code is at: https://github.com/yermandy/GenD.
@@ 1 Introduction
The proliferation of realistic facial deepfakes raises significant concerns regarding misinformation and malicious use, with AI-manipulated videos – those altered by techniques like face swapping or face reenactment – making detection challenging. Unlike fully synthetic content, such forgeries preserve the original context and leave subtle artifacts that are difficult for humans and machines to detect [human-study-1, human-study-2].
@@
The generalization gap is the primary issue that we address in this work. Assuming the hypothesis that adapted large-scale, pre-trained foundational vision encoder can serve as a general foundation for deepfake detection [UniFD], we build the proposed method in three variants, using Contrastive Language-Image Pre-training (CLIP) [CLIP], Perception Encoder (PE) [PE] and DINO [DINOv3] models as feature extractors, which are known for their generalizable visual representations.
The proposed method consists of a vision encoder, whose outputs are L2-normalized. We then fine-tune only the parameters of the Layer Normalization blocks [LN-tuning] while keeping the rest frozen. Additionally, we propose using metric learning in this L2 space to enhance generalization.
@@ 2 Related Work
Deepfake techniques have rapidly evolved. Key methods include: Face-swap [faceswap], Face-reenactment [face-reanectment], LipSync [lipsync], and full face video synthesis [synthesia, veo3]. The first three approaches manipulate only small, localized areas of the video, typically the facial region, while leaving the rest of the footage untouched, making detection challenging. The techniques are continually improving, leveraging advanced models to produce increasingly seamless and realistic results without obvious, visually perceptible artifacts.
As deepfake production techniques advance, detection methods are also evolving to keep pace. The literature encompasses a taxonomy of detection strategies, which can be divided into two main branches: content-based and signal-based approaches. Content-based methods focus on visible or interpretable inconsistencies, utilizing inductive biases such as apparent blending boundaries [SBI], unsynchronized audio-visual movements [LipForensics], or motion artifacts [agarwal2019protecting]. Signal-based methods, on the other hand, detect subtle and often invisible traces left in the visual signal, fingerprints that reveal synthetic origins even when no overt artifacts are present. Signal-based techniques have become increasingly accurate, especially as newer deepfakes leave fewer perceptible cues. However, a persistent challenge remains: generalizing to previously unseen manipulation methods.
Recent efforts in detecting general AI-generated content [UniFD, CLIPping, SSL-in-DF, FatFormer] and facial deepfakes specifically [UDD, LipFD, DFD-FCG, ForensicsAdapter, Effort] have begun to adopt CLIP as a backbone for generalizable deepfake detection. These approaches can be broadly categorized by the type of model adaptation. Some methods, like Forensics Adapter [ForensicsAdapter] (ForAda), introduce architectural changes by adding a separate parallel network explicitly trained to identify artifacts left by blending. Other methods operate on the parameter space; for instance, Effort [Effort] uses Singular Value Decomposition (SVD) to decompose the model’s weights into orthogonal subspaces, freezing the principal components to preserve pre-trained knowledge while fine-tuning the residual components on forgery patterns.
In contrast to these approaches, our work proposes to modify a subset of parameters. LN-Tuning [LN-tuning] adjusts the affine parameters of Layer Normalization blocks. The efficiency of LN-tuning in the general context was analyzed in [LN-tuning-1, LN-tuning-2]. We demonstrate that it achieves competitive or state-of-the-art performance in the deepfake detection task.
As a complimentary approach to the passive detection methods of AI manipulations, there exist active defense approaches that protect real images by incorporating invisible watermarks [zhao2024proactive, asnani2023malp, bartolucci2025perturb, zhai2023defending, asnani2022proactive, yang2021faceguard].
@@ 3.1 Model
Feature extractor. We use three different pre-trained image encoders: 1. CLIP ViT-L/14, trained by OpenAI [CLIP]; 2. $\text{PE}_{\text{core}}\text{L}$, trained by Meta AI [PE]; 3. DINOv3 ViT-L/16 [DINOv3], trained by Meta AI. After the encoder processes the image, we take only the $1024$-dimensional classification token and discard the rest.
Optimized parameters. The L2-normalized classification token is used as the input to the linear binary classifier, which gives logits for fake and real classes. In addition to the weights $W\in\mathbb{R}^{1024\times 2}$ and biases $b\in\mathbb{R}^{2}$ of the classification layer, we optimize the parameters of all Layer Normalization blocks [LN-tuning, LN-tuning-1, LN-tuning-2]. This roughly constitutes 0.03% of the total model parameters; the remaining parameters of the encoder are frozen.
Loss function. The network is trained using a weighted combination of cross-entropy, uniformity, and alignment losses [UniAlign]
@@
Training algorithm. To optimize the model parameters, we use Adam optimizer [Adam] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ without weight decay. The weight precision is set to bfloat16. The learning rate is scheduled using a cosine cyclic rule [cyclic]. Each cycle starts with a linear warm-up for one epoch, increasing from $10^{-5}$ to $3\times 10^{-4}$, and then decays using cosine scheduling over nine epochs to $10^{-5}$. For most runs, we observed that the training did not improve after two cycles (or 20 epochs). The batch size is set to 128 samples. For CLIP ViT-L/14, the loss coefficients of Eq. 1 are set to $\alpha=0.1$ and $\beta=0.5$, found on the validation set.
@@ 3.2 Data
Data preprocessing. We follow a standard practice of dataset preprocessing proposed in DeepfakeBench [DeepfakeBench], which consists of the following steps:
@@
    Extract the largest face using RetinaFace [RetinaFace].
@@
Training data. Following the standard protocol, we train the model on the FaceForensics++ dataset [FF++] (FF++), which consists of $3600$ videos, of which $720$ are real videos, and the rest, $4\times 720$, come from four different deepfake forgery methods. After preprocessing, we obtain approximately 115115k frames. The dataset was released in three different compression levels. Following all of the preceding work, we use the c23 compression version of the dataset. We augment the training set using image augmentations such as random horizontal flipping, random affine transformations, Gaussian blurring, color jitter, and JPEG compression, all implemented in the torchvision package. In ablations, we also experimented with changing the training set to FFIW [FFIW] and DSv2 [DSv1].
@@
Validation data. We noticed that the FF++ validation set is very similar to the training set. That makes the comparison of different training experiments on such a validation ineffective. Each experiment achieves the best validation AUROC quickly, shows almost no signs of overfitting, and does not provide any estimate of cross-dataset generalization. The FF++ validation set does not provide crucial information on how the methods will perform when trained for an extended period. Therefore, we create a custom validation set, which consists of data from validation or training splits of other datasets, namely CDFv3 [CDFv3], DSv1 and DSv2 [DSv1], and FFIW [FFIW]. We ensure that test samples from these datasets are not present in the custom validation set.
@@ 4 Experiments
2019 FF++ [FF++] 4 140 560
2019 DF [FF++] 1 140 140
2019 F2F [FF++] 1 140 140
2019 FS [FF++] 1 140 140
2019 NT [FF++] 1 140 140
2019 DFD [DFD] 5 363 3068 -2
2019 UADFV [UADFV] 1 49 49
2019 DFDC [DFDC] 8 2500 -185 2500 -111
2020 FSh [FSh] 1 140 140
2020 CDFv2 [CDFv2] 1 178 340
2021 FFIW [FFIW] 3 1738 -3 1738 -3
2021 KoDF [KoDF] * 6 403 1106
2021 FAVC [FakeAVCeleb] 4 500 20566 -22
2022 DFDM [DFDM] 5 590 -2 1720 -2
2024 PGF [PolyGlotFake] 10 762 13605
2024 IDF [IDForge] * 9 18834 $23$
2024 DSv1.1 [DSv1] 5 1416 1497
2025 DSv2 [DSv1] 6 1863 1416
2025 CDFv3 [CDFv3] 22 178 2540 -1
@@ 4.1 Test benchmarks
For reporting the performance of the model in cross-dataset settings, we are using Celeb-DF-v2 (CDFv2) [CDFv2], Celeb-DF++ (CDFv3) [CDFv3], DeepFake Detection Challenge (DFDC) [DFDC], Google’s DFD dataset [DFD], Face Forensics in the Wild (FFIW) [FFIW], DeepSpeak v1.1 (DSv1.1) and DeepSpeak v2.0 (DSv2) [DSv1], Korean DeepFake Detection Dataset (KoDF) [KoDF], FakeAVCeleb (FAVC) [FakeAVCeleb], DeepFakes from Different Models (DFDM) [DFDM], PolyGlotFake (PGF) [PolyGlotFake], and IDForge (IDF) [IDForge]. The statistics for the test part of the datasets are shown in  Tab. 1.
@@ 4.2 Cross-dataset evaluation
LipForensics [LipForensics] 2021 CVPR Video ResNet-18 82.4 73.5
@@
RealForensics [RealForensics] 2022 CVPR Video Modified CSN 86.9 75.9
SBI [SBI] 2022 CVPR Frame EFNB4 93.2 82.7 72.4 84.8
AUNet [AUNet] 2023 CVPR Video Xception+ART 92.8 99.2 73.8 81.5
StyleDFD [StyleDFD] 2024 CVPR Video 3D ResNet-50 89.0 96.1
@@
LAA-Net [LAA-Net] 2024 CVPR Frame EFNB4 95.4 98.4 86.9
AltFreezing [AltFreezing] 2024 CVPR Video 3D ResNet-50 89.5 98.5
NACO [NACO] 2024 ECCV Video ViT-B/16 89.5 76.7
RAE [RAE] 2024 ECCV Frame ViT-B/16 95.5 99.0 80.2
TALL++ [TALL++] 2024 IJCV Video Swin-B 92.0 78.5
ProDet [ProDet] 2024 NeurIPS Frame EFNB4 92.5 77.0
UDD [UDD] 2025 AAAI Frame CLIP ViT-B/16 93.1 95.5 81.2
P&P [plug-and-play] 2025 CVPR Video CLIP ViT-L/14 94.7 96.5 84.3 92.1
DFD-FCG [DFD-FCG] 2025 CVPR Video CLIP ViT-L/14 95.0 81.8
ForAda [ForensicsAdapter] 2025 CVPR Frame CLIP ViT-L/14 95.7 97.2 87.2
Effort [Effort] 2025 ICML Frame CLIP ViT-L/14 95.6 96.5 84.3 92.1
@@
The proposed method achieves these results simply by averaging individual frame softmax probabilities, without complex post-processing or architectural modifications as described in Sec. 3.1. This contrasts with many contemporary methods that rely on complex multimodal analyses, such as modeling temporal inconsistencies between video frames or detecting audio-video discrepancies [LipForensics, LipFD, RealForensics, FTCN, DFD-FCG, plug-and-play]. The performance of the proposed method suggests that the features of the visual encoder, when properly tuned, are highly effective for this detection task.
We also evaluated the generalization of the proposed method against current SOTA models across a wide range of benchmark datasets. For this experiment, we selected the most recent open-source SOTA models that have made their code and pre-trained weights publicly available, namely ForAda [ForensicsAdapter] and Effort [Effort]. To ensure that each model was used as intended by its original authors, we utilized their officially released weights and integrated their data preprocessing pipelines which were used during model training. After reproducing the results on the test datasets that were used in these papers, we obtained the same or very similar results, presented in Tab. 5. The in-distribution results on FF++ are shown in Tab. 3.
@@
ForAda [ForensicsAdapter] 99.7 97.0 98.6 91.9 96.8
Effort [Effort] 99.4 93.2 98.4 84.6 93.9
@@
ForAda [ForensicsAdapter] 99.4 97.2 87.3 82.0 95.7 90.6 88.2 93.1 97.1 86.6 90.8 81.8 72.8 75.6 88.4
Effort [Effort] 97.4 95.2 85.4 91.2 93.2 92.5 88.1 92.4 98.2 84.9 96.0 82.1 64.4 78.7 88.5
@@
ForAda [ForensicsAdapter] 95.7 / 95.7 97.2 / 97.2 87.2 / 87.3 – / 90.6
Effort [Effort] 95.6 / 93.2 96.5 / 95.2 84.3 / 85.4 92.1 / 92.5
@@ 4.3 Ablation studies
    Baseline: the setup is similar to [UniFD]. We freeze the backbone and train a linear classifier with cross-entropy loss on top of the feature space of the classification token.
@@
We explore further parameter-efficient fine-tuning (PEFT) strategies, the goal of which is to modify only a relatively small number of parameters while keeping most parameters untouched, thus efficiently introducing several degrees of freedom. PEFT is more effective for generalization than full fine-tuning (FFT) in low-data regimes [PEFT], which is particularly applicable to deepfake detection, where collected datasets and benchmarks have a limited number of samples. This work explores three PEFT methods: LoRA [hu2022lora], LN-tuning [LN-tuning, CLIPFit], and bias tuning (BitFit) [CLIPFit, BitFit].
We found that FFT leads to rapid overfitting and poor generalization across datasets. This observation is consistent with recent findings from Effort [Effort] that FFT degrades the generalization performance.
Similarly to [Effort], we fine-tuned the parameters of every transformer weight matrix using LoRA to preserve the generalization. Setting LoRA to rank one, the lowest possible, allowed the model to reach a near-perfect training AUROC (99.99%) in two epochs, resulting in rapid overfitting.
We observed that BitFit [BitFit] and LN-tuning [LN-tuning] have stable training dynamics by first improving the validation AUROC, reaching a peak, and overfitting after some epochs. In comparison, FFT and LoRA show no improvement in validation AUROC even after a single training epoch.
@@
Uniformity and Alignment. We hypothesized that representations that effectively exploit the latent space could help us improve generalization. We used uniformity and alignment loss [UniAlign], leading to slower overfitting and better validation AUROC. We suppose that adding these losses helps to utilize the hyperspherical space more effectively, resulting in improved generalization.